# A robust phase of continuous transversal gates in quantum stabilizer codes

### v1

@@ Abstract
The conventional strategy for fault-tolerant non-transversal gates is to produce magic states for injection by gate teleportation [2]. This requires mid-circuit measurement and feed-forward, which has been demonstrated experimentally in linear optics [3, 4, 5], trapped ions [6, 7], superconducting circuits [8, 9], and neutral atoms [10, 11, 12]. Magic state distillation produces high fidelity magic states from many lower quality ones [13]. Numerous refinements have reduced its resource requirements [14, 15, 16, 17, 18, 19], while alternative methods of growing magic states with syndrome extraction and postselection  [20, 21, 22] have culminated in magic state cultivation [23, 24, 25, 26].
Although distillation schemes may be optimized for small rotation angles [27], most protocols produce fixed-angle rotations so synthesizing arbitrary angle rotations up to some tolerance $\epsilon$ still requires $O(\log(1/\epsilon))$ operations [28]. This becomes costly for small rotations where we want $\epsilon$ to be much smaller than the rotation. Thus non-distillation schemes for continuous angles based on rotations and measurements with postselection have gained interest [29, 30]. Practical limitations of near-term devices have also motivated partially fault-tolerant analog rotations [31, 32].
@@
When this logical rotation is repeated for $t$ rounds, as shown in Fig. 1(b), with optimally chosen rotation angles $\theta_{j}$ for each round $j$, then the logical rotation angles will sum to a total logical rotation angle $\Phi_{t}$ which can be steered towards to a target a non-Clifford logical gate.
It thus remains to show the fault-tolerance of the protocol. $X$ errors on the input state or before the first round of $X$ syndrome measurements merely cause logical $X$ errors that do not affect the logical $Z$ angle $\phi_{s}$ due to the identity in Eq. (1). The protocol can also be made fault-tolerant against measurement faults and Pauli errors that occur after the first round of syndrome measurement. To see this note that the syndrome extraction projects the state into a syndrome subspace, fixing the coherent rotation angle up to a Pauli error, so to be robust against these errors we just need to determine which syndrome subspace the system was in at the first round. This initial value of the syndrome can be learned by using $O(d)$ rounds of conventional syndrome extraction (see Appendix)  [46], or by other means [56]. On the other hand, since $Z$ errors on the input state, or following the coherent rotation, cause syndromes from the first round of syndrome extraction that are indistinguishable from coherent $Z$ rotations, the applied logical rotation angle is no longer knowable with certainty in the presence of $Z$ noise. The effect of dephasing hence warrants a more careful study, which is the focus of this work.
@@
The $\theta=0$ case corresponds to dephasing only, with the logical rotation angle $\phi_{s}=0$ for all syndromes $s$. In the low-$p$ and small $\theta$ regime, the most common syndrome is the trivial syndrome $\vec{0}$ so the $\phi_{s}$ distribution is peaked the angle $\phi_{0}$ corresponding to the trivial syndrome. Above the error correction threshold, $\phi_{s}$ becomes almost uniformly distributed. More properties of this distribution are listed in the Appendix.
@@
Figure 2: (a) Phase diagram in physical chanel parameter space $(p,\theta)$. In the robust logical coherent phase (shaded pink), the mean relative dephasing $\mathbb{E}[q_{s}/|\phi_{s}|]\to 0$ as $d\to\infty$ using the PyMatching decoder [58]. It is contained within the correctable phase for conventional QEC storage (shaded green) where $\left\|\overline{\mathcal{E}}\ -\mathcal{I}\right\|_{\diamond}\to 0$ as $d\to\infty$. (b) For a fixed physical dephasing $p=0.001$, the mean relative dephasing has a minimum value at a $\theta$ in the robust phase that is bounded from above and below by phase transitions. (c) The logical relative dephasing at the target angle that succeeds with 50% probability is exponentially suppressed with the distance of the code.
@@
For sufficiently small $\Phi_{T}$, the probability of success in the first round is the probability $p(\vec{0})$ of obtaining the trivial syndrome $s=\vec{0}$ with $\theta$ chosen such that $\phi_{0}=\Phi_{T}$. Fig. 2(c) shows that the resulting relative dephasing $q_{0}/|\phi_{0}|$ is suppressed exponentially with distance when exactly targeting the logical angle $\Phi_{T}$ at which $p(\vec{0})\approx 50\%$.
@@
Figure 3: Performance of the adaptive protocol with resets for preparing magic states with logical target angle $\Phi_{T}$ at physical dephasing rate $p=0.001$. (a) The expected number of rounds $\mathbb{E}[T]$ including resets is independent of the tolerance $\epsilon$, but increases with target angle $\Phi_{T}$. (b) The total logical dephasing $\mathbb{E}[Q_{T}]$ increases with the target angle $\Phi_{T}$. (c) There is a target angle where the relative logical dephasing $\mathbb{E}[Q_{T}/|\Phi_{T}|]$ is minimal. The dashed lines indicate the target $\Phi_{T}$ at which the protocol will succeed in one just round with 50% probability with the color corresponding to the code distance.
@@
Optimizing for $\Phi_{t}$ to converge to a target $\Phi_{T}$ without regards to $Q_{t}$ will result in excessively high $Q_{t}$ values. If our goal is to prepare a magic state for injection, we have the option to reset and start afresh. We can generalize the state space to $(\Phi,Q)$ and optimize the policy similarly to minimize the total number rounds including resets. Details of the policy optimization and simulation with dephasing are in the Appendix. The expected number of rounds $\mathbb{E}[T]$ needed to achieve the target angle $\Phi_{T}$ estimated using this method is plotted in Fig. 3(a). The resulting mean total dephasing $\mathbb{E}[Q_{T}]$ depends on the physical dephasing rate $p$ and target logical angle $\Phi_{T}$. This total dephasing rate is plotted for various target angles in Fig. 3(b). We observe in Fig. 3(c) that as the code distance $d$ increases, the mean logical relative dephasing $\mathbb{E}[Q_{T}/|\Phi_{T}|]$ is minimized near the $\Phi_{T}$ at which one round would succeed with 50% probability.
Discussion—In this Letter, we have shown the existence of a robust phase of continuous transversal gates in the surface code. We exploit this to introduce a low-cost adaptive protocol that efficiently implements continuously tunable fault-tolerant non-Clifford gates in the surface code over a range of rotation angles using only transversal operations and syndrome measurements. The rotation is exact and postselection was avoided with subsequent rotations and resets when preparing resource states for injection. This protocol presents a powerful new addition to the toolkit for fault-tolerant non-Clifford gates.
@@
A natural extension of this work is to assess performance against phenomenological and circuit-level noise models using the full syndrome extraction circuits, for which we expect the results to hold qualitatively by leveraging existing fault-tolerant protocols of the surface code. It is expected that the protocol should apply for odd-distance codes satisfying the logical unitary criterion in [55], which may include more efficient quantum low-density parity check (LDPC) codes [74, 75, 76, 77]. Further gains may be obtained by non-uniform physical qubit rotations and better decoders. Experimental demonstration of our protocol will be a future challenge.

### v2

@@ Abstract
The conventional strategy for fault-tolerant non-transversal gates is to produce magic states for injection by gate teleportation [2]. This requires mid-circuit measurement and feed-forward, which has been demonstrated experimentally in linear optics [3, 4, 5], trapped ions [6, 7], superconducting circuits [8, 9], and neutral atoms [10, 11, 12]. Magic state distillation produces high fidelity magic states from many lower quality ones [13]. Numerous refinements have reduced its resource requirements [14, 15, 16, 17, 18, 19], while alternative methods of growing magic states with syndrome extraction and postselection  [20, 21, 22] have culminated in magic state cultivation [23, 24, 25, 26, 27].
Although distillation schemes may be optimized for small rotation angles [28], most protocols produce fixed-angle rotations so synthesizing arbitrary angle rotations up to some tolerance $\epsilon$ still requires $O(\log(1/\epsilon))$ operations [29]. This becomes costly for small rotations where we want $\epsilon$ to be much smaller than the rotation. Thus non-distillation schemes for continuous angles based on rotations and measurements with postselection have gained interest [30, 31]. Practical limitations of near-term devices have also motivated partially fault-tolerant analog rotations [32, 33, 34, 35].
@@
When this logical rotation is repeated for $t$ rounds, as shown in Fig. 1(b), with optimally chosen rotation angles $\theta_{j}$ for each round $j$, then the logical rotation angles will sum to a total logical rotation angle $\Phi_{t}$ which can be steered towards to a non-Clifford target logical angle $\Phi_{T}$.
It thus remains to show the fault-tolerance of the protocol. $X$ errors on the input state or before the first round of $X$ syndrome measurements merely cause logical $X$ errors that do not affect the logical $Z$ angle $\phi_{s}$ due to the identity in Eq. (1). The protocol can also be made fault-tolerant against measurement faults and Pauli errors that occur after the first round of syndrome measurement. To see this note that the syndrome extraction projects the state into a syndrome subspace, fixing the coherent rotation angle up to a Pauli error, so to be robust against these errors we just need to determine which syndrome subspace the system was in at the first round. This initial value of the syndrome can be learned by using $O(d)$ rounds of conventional syndrome extraction (see Appendix)  [49], or by other means [59]. On the other hand, since $Z$ errors on the input state, or following the coherent rotation [60], cause syndromes from the first round of syndrome extraction that are indistinguishable from coherent $Z$ rotations, the applied logical rotation angle is no longer knowable with certainty in the presence of $Z$ noise. The effect of dephasing hence warrants a more careful study, which is the focus of this work.
@@
The $\theta=0$ case corresponds to dephasing only, with the logical rotation angle $\phi_{s}=0$ for all syndromes $s$. In the low-$p$ and small $\theta$ regime, the most common syndrome is the trivial syndrome $\vec{0}$ so the $\phi_{s}$ distribution is peaked at the angle $\phi_{0}$ corresponding to the trivial syndrome. Above the error correction threshold, $\phi_{s}$ becomes almost uniformly distributed. More properties of this distribution are listed in the Appendix.
@@
Figure 2: (a) Phase diagram in physical chanel parameter space $(p,\theta)$. In the robust logical coherent phase (shaded pink), the mean relative dephasing $\mathbb{E}[q_{s}/|\phi_{s}|]\to 0$ as $d\to\infty$ using the PyMatching decoder [62]. It is contained within the correctable phase for conventional QEC storage (shaded green) where the diamond distance from the identity channel $\left\|\overline{\mathcal{E}}\ -\mathcal{I}\right\|_{\diamond}\to 0$ as $d\to\infty$. (b) For a fixed physical dephasing $p=0.001$, the mean relative dephasing has a minimum value at a $\theta$ in the robust phase that is bounded from above and below by phase transitions. (c) The logical relative dephasing at the target angle that succeeds with 50% probability is exponentially suppressed with the distance of the code.
@@
For sufficiently small target logical angle $\Phi_{T}$, the probability of success in the first round is the probability $p(\vec{0})$ of obtaining the trivial syndrome $s=\vec{0}$ with $\theta$ chosen such that $\phi_{0}=\Phi_{T}$. Fig. 2(c) shows that the resulting relative dephasing $q_{0}/|\phi_{0}|$ is suppressed exponentially with distance when exactly targeting the logical angle $\Phi_{T}$ at which $p(\vec{0})\approx 50\%$.
@@
Figure 3: Performance of the adaptive protocol with resets for preparing magic states with logical target angle $\Phi_{T}$ at physical dephasing rate $p=0.001$. (a) The expected number of rounds $\mathbb{E}[T]$ including resets is independent of the tolerance $\epsilon$, but increases with target angle $\Phi_{T}$. (b) There is a target angle where the relative logical dephasing $\mathbb{E}[Q_{T}/|\Phi_{T}|]$ is minimal. The dashed lines indicate the target $\Phi_{T}$ at which the protocol will succeed in just one round with 50% probability with the color corresponding to the code distance. (c) Resource comparison of adaptive protocol against the limits of magic state cultivation from Gidney et al. [23] with Clifford+T synthesis. Each point denotes a target angle $\Phi_{T}$. Details are in the Appendix.
@@
Optimizing for $\Phi_{t}$ to converge to a target $\Phi_{T}$ without regards to $Q_{t}$ will result in excessively high $Q_{t}$ values. If our goal is to prepare a magic state for injection, we have the option to reset and start afresh. We can generalize the state space to $(\Phi,Q)$ and optimize the policy similarly to minimize the total number rounds including resets. Details of the policy optimization and simulation with dephasing are in the Appendix. The expected number of rounds $\mathbb{E}[T]$ needed to achieve the target angle $\Phi_{T}$ estimated using this method is plotted in Fig. 3(a). The mean logical relative dephasing $\mathbb{E}[Q_{T}/|\phi_{T}|]$ depends on the physical dephasing rate $p$ and target logical angle $\Phi_{T}$, as shown in Fig. 3(b). As the code distance $d$ increases, its value is minimized near the target angle $\Phi_{T}$ at which one round would succeed with 50% probability.
Discussion—In this Letter, we have shown the existence of a robust phase of continuous transversal gates in the surface code. We exploit this to introduce a low-cost adaptive protocol that efficiently implements continuously tunable fault-tolerant non-Clifford gates in the surface code over a range of rotation angles using only transversal operations and syndrome measurements. The rotation is exact and postselection was avoided with subsequent rotations and resets when preparing resource states for injection. This protocol presents a powerful new addition to the toolkit for fault-tolerant non-Clifford gates. Its estimated total logical error rate and cost in terms of expected spacetime volume including resets for various target angles and distances is plotted in Fig. 3(c), where the achievable limits of conventional magic state cultivation and Clifford+T synthesis are also included for reference. For example, the $d=7$ protocol becomes competitive when targeting $\Phi_{T}\leq\pi/2^{8}$.
@@
A natural extension of this work is to assess performance under phenomenological and circuit-level noise models with full syndrome extraction circuits, for which we expect the results to hold qualitatively by leveraging existing fault-tolerant protocols of the surface code. It is expected that the protocol should apply for odd-distance codes satisfying the logical unitary criterion in [58], which may include more efficient quantum low-density parity check (LDPC) codes [78, 79, 80, 81]. Further gains may be seen by non-uniform physical qubit rotations, two-qubit rotations [35], and better decoders. Experimental demonstration of our protocol will be a future challenge.